# Identifying Important Group of Pixels using Interactions

### v1

@@ 4.2 Pixel Deletion
In Problem 1, we solved the issue of identifying groups of pixels with high confidence scores when additional pixels are added. In pixel deletion, we address the problem of identifying groups of pixels that result in low confidence scores when they are removed.
@@
With these quantities, the greedy algorithm for pixel deletion is as follows. For $k=1$, the index $b_{1}\in N$ of the pixel with the lowest (deletion-based) Shapley value $-\phi_{\mathrm{d}}^{(|N|)}(b_{1})(=\frac{1}{|N|}\>[f(N\setminus\{a\})-f(N)])$ gives the optimal set $S_{1}=\{b_{1}\}$ by its definition. For $k=2$, we select the next pixel $b_{2}$ that minimizes $f(N\setminus\{b_{1},b_{2}\})$. This choice is again explained as a sum of Shapley value and interaction,
@@
$=\underset{b\in N\setminus\{b_{1}\}}{\mathrm{arg\,max}}\ \phi_{\mathrm{d}}^{(|N|)}(b)+(|N|-1)I_{d}^{(|N|-1)}(b_{1},b).$ (12)
@@
$\quad+(|N|-|S_{k-1}|)\,I_{d}^{(|N|-|S_{k-1}|)}(S_{k-1},b)].$ (13)
Again, the greedy algorithm is described from a game-theoretic viewpoint. The only difference from the insertion case is that the interaction term is now weighted. Algorithm 2 summarises the procedure.
The computational cost of the pixel deletion approach is the same as the pixel insertion approach, which only requires $\mathcal{O}(|N|)$ times of forward passes.
@@ 5 Experiments
Our experiments utilize the ImageNet dataset [10] and focus on analyzing Vision Transformer [11] pre-trained for the classification task. For baseline methods, we use Grad-CAM [23] [^2] The target layer of Grad-CAM is set to the one before the layer normalization in the final attention block of network. This choice is common, see https://github.com/jacobgil/pytorch-grad-cam., Attention rollout [1], Shapley values, and MoXI(-), which do not utilize the interactions present in MoXI. For insertion curve experiments, we use the Pixel Insertion approach with MoXI(-) and MoXI, while for deletion curves, we utilize the Pixel Deletion approach with MoXI(-) and MoXI. Following the previous studies [21, 29], we consider image patches instead of pixels to reduce computational costs in the experiments. All methods calculate the contributions for $14\times 14$ patches with a patch size of $16\times 16$, which is equal to the patch size and the number of tokens in standard ViT models. We used a pre-trained ViT-T [11] [^3] https://huggingface.co/WinKawaks/vit-tiny-patch16-224 and DeiT-T [26] [^4] https://huggingface.co/facebook/deit-tiny-patch16-224. We selected $1000$ images, one corresponding to each label, all of which were successfully classified in the test set. To reduce the computational burden, we computed Shapley values approximately by random sampling of $S$ in Eq. (1) as in other studies [4, 22, 29, 25]. The sampling size is set to 200. Additionally, we have adopted feature patch deletion as the masking method for Shapley values and interactions. In the following, we focus on ViT-T. See Appendix B for more results.
@@ 5.1 Evaluating the importance of identified patches
Figure 3: (a) Insertion curves. (b) Deletion curves. The curves illustrate the change of accuracy when adding (removing) image patches gradually with high contributions identified by different methods at various unmasked (masked) image rates, ranging from 00 to $100\%$. The masking method used in game-theoretic computations employs patch deletion. For the insertion/deletion curve experiments, masks used for input to the model for accuracy measurement also employ patch deletion.
@@
The insertion curves in Fig. 3(a) show that MoXI exhibits a sharper increase in classification accuracy compared to the other methods. In particular, even with images where only $4\%$ is visible, MoXI achieves an accuracy of $90\%$, whereas Grad-CAM, Attention rollout, and Shapley value achieve $2\%$, $4\%$, and $25\%$, respectively. This result indicates that MoXI can efficiently identify patches important in the model’s decision-making process. Then, both the self-context and original Shapley values, which are based on confidence scores, achieve a sharper increase in classification accuracy. However, these two methods calculate the importance of individual patches and often select patches with similar information. Consequently, MoXI can identify features contributing to a higher classification accuracy than these methods.
The deletion curves in Fig. 3(b) show that MoXI exhibits a sharp decrease in classification accuracy compared to the other methods. When concealing just $10\%$ of an image, MoXI significantly decreases the model’s accuracy to $16\%$. In contrast, Grad-CAM and Attention rollout only decrease the accuracy to approximately $79\%$ under the same conditions. This result indicates that MoXI, which accounts for interactions between patches, effectively identifies the image patches important for classification. We observed analogous results for DeiT-T [26] models, as detailed in Appendix B. Additionally, we discuss the application of masks using our proposed method in Appendix C.
@@ 5.2 Confidence score-based visualization
Figure 4(a) displays a heatmap for patch insertion. Compared to the existing methods, MoXI’s heatmap highlights fewer regions and identifies the class object. Interestingly, MoXI selects the patches on the background as well as the class object. This visualization explains the object and background is required for classification and demonstrates the usefulness of the interaction. Figure 4(b) displays a heatmap for patch deletion. The heatmaps generated by MoXI(-) and Grad-CAM display extensive highlights across the image, while MoXI, Attention rollout, and Shapley value show more concentrated highlights on the class object. This finding indicates that these latter methods accurately capture important information from the object. Notably, MoXI places less emphasis on the background than Attention rollout and Shapley value. This result suggests that MoXI effectively narrows down information by selectively deleting the class object, which could be advantageous for precise object localization.
@@
To enhance understanding of the model’s prediction process, localization for specific classes improve interpretability. We have extended MoXI to analyze a target class that differs from the model’s prediction. For the detailed of visualization, see Appendix E. Figures 5(b) and (c) visualizes important regions for two classes: the bull mastiff, as predicted by the model, and the tiger cat, the target class. The heatmaps reveal that MoXI highlights the bull mastiff’s facial area and the tiger cat’s face and body. These observations demonstrate that MoXI can identify important groups of image patches relevant to the predicted class and class-specific features important for decision-making.
@@ 5.3 Common corruption effect on patch deletion
We investigate the risk of model misclassification when image patches important for model accuracy are disrupted by adding noise. In the deletion curve experiment of Sec. 5.1, we used patch masking to simulate feature absence. Instead of patch masking, we consider common corruption [15]: fog and Gaussian noise at level 5 (for the other corruptions such as brightness and motion blur, see Appendix F.1). We apply these corruptions to image patches in the order selected for patch deletion in Sec. 5.1. Figure 6(a) shows the effect of Gaussian noise on the deletion curve results. MoXI exhibits a significant decrease in accuracy compared to the others, indicating MoXI is vulnerable to Gaussian noise. This result implies that MoXI efficiently identifies important patches for classification. Figure 6(b) shows the fog corruption results, which are similar to those observed for Gaussian noise. Furthermore, as detailed in Appendix F.1, MoXI similarly affects accuracy with the other common corruptions. Additionally, we evaluate the effect of adversarial perturbations. Interestingly, adversarial perturbations yield distinct results due to their deceptive effect on the model’s internal features (see Appendix F.2).
Figure 6: Deletion curves. The curves illustrate the change in accuracy when gradually applying noise to image patches with high contributions identified by different methods at various noise-added image rates, ranging from 00 to $100\%$. The order of applying noise to image patches is based on the sequence obtained from the deletion curve experiment in Fig. 5. The types of noise applied are (a) Gaussian noise and (b) fog.
@@ 5.4 Consistant explainability
We examine the consistent explainability of the visualization methods, irrespective of the internal feature representation, which is a key aspect of explainable artificial intelligence. Specifically, we examine whether the models, trained with varying numbers of classification classes, consistently select important image patches. We evaluate the consistency using insertion and deletion curves for the models trained with datasets containing 10, 20, 100, and 1000 classes. For training the 10-class model, we select images from ImageNet that share labels with CIFAR10. For the models with 20, 100, and 1000 classes, we extend the 10-class dataset by adding images with randomly selected classes from ImageNet. We draw the insertion and deletion curves using the 10-class test images that are correctly classified. Figures 7(a) and (b) shows the insertion curve results for Attention rollout and MoXI, respectively. Attention rollout decreases accuracy as the number of classes increases. In contrast, MoXI does not decrease in accuracy. Therefore, MoXI consistently selects important image patches for accurate classification. In addition, the results from other methods and deletion experiments are shown in the Appendix G. We confirmed that MoXI provides consistent explainability in the deletion curve experiments.
Figure 7: Insertion curves. The curves illustrate the change in accuracy when gradually applying noise to image patches with high contributions identified by different methods at various noise-added image rates, ranging from 00 to $100\%$. (a) Results for Attention rollout, (b) Results for MoXI.
@@ 6 Conclusion
This study addressed the problem of identifying a group of pixels that largely and collectively impact confidence scores in image classification models. We justify simple greedy algorithms from a game-theoretic view using Shapley values and interactions. This analysis naturally suggests the use of self-context and full-context variants of Shapley values and interactions. Their computation only requires a linear number of forward passes, whereas prior studies compute Shapley values and/or interactions with an exponential number of forward passes or heavy sampling-based approximation. The experimental results show that our method is more accurate in identifying the important image patches for models than popular methods such as Grad-CAM, Attention rollout, and Shapley value.
@@ A Visual Set-Sum Task
B Insertion and deletion curves for DeiT-T
@@
In Sec. 5.1, we evaluated the proposed and baseline methods using ViT-T [11]. The insertion and deletion curves show that the proposed method provides the most efficient visual explanation. To show this generalization to other models, we provide the results with DeiT-T [26], a model extensively utilized in ViT architectures. For details of the experiment, refer to Sec. 5.1. The insertion curve in Fig. 8(a) again shows that MoXI exhibits a sharper increase compared to the other methods. Similarly, the deletion curve in Fig. 8(b) demonstrated that MoXI exhibits a shaper decrease compared to the other methods. These results indicate that our method in DeiT-T can efficiently and accurately identify the critical patches in the model’s decision-making process.
@@
Figure 8: (a) Insertion curves. (b) Deletion curves. The curves illustrate the accuracy growth when inserting (deleting) image patches according to the contributions computed by each method. The horizontal axis presents the insertion rate.
C Analysis of effective layers to remove patches
In Sec. 5.1, we consider the absence of players (i.e., pixels/patches) for calculating Shapley values and interactions in the input space. Specifically, the patches are removed after the input embedding layer.
Here, we examine the case where several self-attention layers are instead masked. To this end, we utilize a variant of the attention-masking approach used in [8].
@@ C Analysis of effective layers to remove patches
Specifically, let the $k$-th layer be our target layer. Then, a large negative value is added to the product of the query and key matrices from $k$-th to the last self-attention layers. Figure 9 displays the insertion curve results when MoXI is applied to various target layers. The experimental setup is the same as in Sec. 5.1. The result demonstrates that MoXI prefers the earlier layers and better pinpoints the important features of images.
@@
D Additional results of visualization
We provide additional visualization results in Fig. 10. As in Sec. 5.2, the results demonstrated that the patches highlighted by MoXI are smaller than those highlighted by other methods.

### v2

@@ 4.2 Pixel Deletion
To address Problem 1, we considered the problem of identifying groups of pixels with high confidence scores through pixel insertion. Here, we aim at decreasing the confidence scores via pixel deletion.
@@
With these quantities, the greedy algorithm for pixel deletion is as follows. For $k=1$, the index $b_{1}\in N$ of the pixel with the highest (deletion-based) Shapley value $\phi_{\mathrm{d}}^{(|N|)}(b_{1})=-\frac{1}{|N|}\>[f(N\setminus\{a\})-f(N)]$ gives the optimal set $S_{1}=\{b_{1}\}$ by its definition. For $k=2$, we select the next pixel $b_{2}$ that minimizes $f(N\setminus\{b_{1},b_{2}\})$. This choice is again explained as a sum of Shapley value and interaction,
@@
$=\underset{b\in N\setminus\{b_{1}\}}{\mathrm{arg\,max}}\ \phi_{\mathrm{d}}^{(|N|)}(b)+(|N|-1)I_{\mathrm{d}}^{(|N|-1)}(b_{1},b).$ (12)
@@
$\quad+(|N|-|S_{k-1}|)\,I_{\mathrm{d}}^{(|N|-|S_{k-1}|)}(S_{k-1},b)].$ (13)
@@
Again, the greedy algorithm is described from a game-theoretic viewpoint. The only difference from the insertion case is that the interaction term is now weighted. Algorithm 2 summarises the procedure. The computational cost of the pixel deletion approach is the same as the pixel insertion approach, which only requires $\mathcal{O}(|N|^{2})$ times of forward passes in the worst case.
@@ 5 Experiments
Our experiments utilize the ImageNet dataset [10] and focus on analyzing Vision Transformer [11] pre-trained for the classification task. For baseline methods, we use Grad-CAM [23] [^2] The target layer of Grad-CAM is set to the one before the layer normalization in the final attention block of network. This choice is common, see https://github.com/jacobgil/pytorch-grad-cam., Grad-CAM++ [5], Attention rollout [1], Shapley values, and MoXI(-), which do not utilize the interactions present in MoXI. For insertion curve experiments, we use the Pixel Insertion approach, while for deletion curves, we utilize the Pixel Deletion approach. Following the previous studies [21, 29], we consider image patches instead of pixels to reduce computational costs. All methods calculate the contributions for $14\times 14$ patches with a patch size of $16\times 16$, which is equal to the patch size and the number of tokens in standard ViT models. We used a pre-trained ViT-T [^3] https://huggingface.co/WinKawaks/vit-tiny-patch16-224 [11], DeiT-T [^4] https://huggingface.co/facebook/deit-tiny-patch16-224 [26] and ResNet-18555https://huggingface.co/microsoft/resnet-18 [14]. We selected $1000$ images, one corresponding to each label, all of which were successfully classified in the test set. To reduce the computational burden, we computed Shapley values approximately by random sampling of $S$ in Eq. (1) as in other studies [4, 22, 29, 25]. The sampling size is set to 200. Moreover, we have adopted feature patch deletion as the masking method for Shapley values and interactions. In the following, we focus on ViT-T. See Appendix B for more results.
@@ 5.1 Evaluating the importance of identified patches
Figure 3: (a) Insertion curves. (b) Deletion curves. The curves illustrate the change of accuracy when appending (removing) image patches gradually with high contributions identified by different methods at various unmasked (masked) image rates, ranging from 00 to $100\%$.
@@
The insertion curves in Fig. 3(a) show that MoXI exhibits a sharper increase in classification accuracy compared to the other methods. In particular, even with images where only $4\%$ is visible, MoXI achieves an accuracy of $90\%$, whereas Grad-CAM, Attention rollout, and Shapley value achieve $2\%$, $4\%$, and $25\%$, respectively. This result indicates that MoXI can efficiently identify important patches for classification. Then, both the self-context and original Shapley values, which are based on confidence scores, achieve a sharper increase in classification accuracy. However, these two methods calculate the importance of individual patches and often select patches with similar information. Consequently, MoXI can identify features contributing to a higher classification accuracy than these methods.
The deletion curves in Fig. 3(b) show that MoXI exhibits a sharp decrease in classification accuracy compared to the other methods. When concealing just $10\%$ of an image, MoXI significantly decreases the model’s accuracy to $16\%$. In contrast, Grad-CAM and Attention rollout only decrease the accuracy to approximately $79\%$ under the same conditions. This result indicates that MoXI, which accounts for interactions between patches, effectively identifies the image patches important for classification. We observed analogous results for DeiT-T [26] and ResNet-18 [14] models, as detailed in Appendix B. Additionally, we discuss the application of masks using our method in Appendix D.
@@ 5.2 Confidence score-based visualization
Figure 4(a) displays a heatmap for patch insertion. Compared to the existing methods, MoXI’s heatmap highlights fewer regions and identifies the class object. Interestingly, MoXI selects the patches on the background as well as the class object. This visualization explains the object and background is required for classification and demonstrates the usefulness of the interaction.
Figure 4(b) displays a heatmap for patch deletion. The heatmaps generated by MoXI(-) and Grad-CAM display extensive highlights across the image, while MoXI, Attention rollout, and Shapley value show more concentrated highlights on the class object. This finding indicates that these latter methods accurately capture important information from the object. Notably, MoXI places less emphasis on the background than Attention rollout and Shapley value. This result suggests that MoXI effectively narrows down information by selectively deleting the class object, which could be advantageous for precise object localization.
@@
To enhance understanding of the model’s prediction process, localization for specific classes improve interpretability. We have extended MoXI to analyze a target class that differs from the model’s prediction. For the detailed visualization, see Appendix F. Figures 5(b) and  5(c) visualizes important regions for two classes: the bull mastiff, as predicted by the model, and the tiger cat, the target class. The heatmaps reveal that MoXI highlights the bull mastiff’s facial area and the tiger cat’s face and body. These observations demonstrate that MoXI can identify important groups of image patches relevant to the predicted class and class-specific features important for decision-making.
@@ 5.3 Common corruption effect on patch deletion
We investigate the risk of model misclassification when image patches important for model accuracy are disrupted by adding noise. In the deletion curve experiment of Sec. 5.1, we used patch masking to simulate feature absence. Instead of patch masking, we consider common corruption [15]: fog and Gaussian noise at level 5 (for the other corruptions such as brightness and motion blur, see Appendix G.1). We apply these corruptions to image patches in the order selected for patch deletion in Sec. 5.1.
Figure 6(a) shows the effect of Gaussian noise on the deletion curve results. MoXI exhibits a significant decrease in accuracy compared to the others, indicating MoXI is vulnerable to Gaussian noise. This result implies that MoXI efficiently identifies important patches. Figure 6(b) shows the fog corruption results, which are similar to those observed for Gaussian noise. Furthermore, as detailed in Appendix G.1, MoXI similarly affects accuracy with the other common corruptions. Additionally, we evaluate the effect of adversarial perturbations. Interestingly, adversarial perturbations yield distinct results due to their deceptive effect on the model’s internal features (see Appendix G.2).
Figure 6: Deletion curves by image corruptions instead of masking: (a) Gaussian noise and (b) fog. The curves illustrate the change in accuracy along with the increase in the number of corrupted image patches. The patches are corrupted from the highly contributing ones determined by each method.
@@ 5.4 Consistant explainability
We examine the consistent explainability of visualization methods, regardless of the internal feature representation, which is a key aspect of explainable artificial intelligence. Specifically, we examine whether the models, trained with varying numbers of classification classes, consistently select important image patches. We evaluate the consistency using insertion and deletion curves for the models trained with datasets containing 10, 20, 100, and 1000 classes. For training the 10-class model, we select images from ImageNet that share labels with CIFAR10. For the models with 20, 100, and 1000 classes, we extend the 10-class dataset by adding images with randomly selected classes from ImageNet. We draw the insertion and deletion curves using the 10-class test images that are correctly classified.
Figures 7(a) and 7(b) show the insertion curve results for Attention rollout and MoXI, respectively. Attention rollout decreases accuracy as the number of classes increases. In contrast, MoXI does not decrease in accuracy. Therefore, MoXI consistently selects important image patches for accurate classification. In addition, the results from other methods and deletion experiments are shown in the Appendix H. We confirmed that MoXI provides consistent explainability in the deletion curve experiments.
Figure 7: Insertion curves. (a) Attention rollout, (b) MoXI. The curves illustrate the change in accuracy along with the increase in the number of unmasked image patches. Each curve represents the results from the pretrained models with $10$, $20$, $100$, and $1000$ classes, respectively. As the number of classes the model learns increases, the accuracy of Attention rollout significantly decreases, whereas MoXI experiences only a minor decrease in accuracy.
@@ 6 Conclusion
This study addressed the problem of identifying a group of pixels that largely and collectively impact confidence scores in image classification models. We justify simple greedy algorithms from a game-theoretic view using Shapley values and interactions. This analysis naturally suggests the use of self-context and full-context variants of Shapley values and interactions. Their computation only requires a quadratic number of forward passes, whereas prior studies compute Shapley values and/or interactions with an exponential number of forward passes or heavy sampling-based approximation. The experimental results show that our method is more accurate in identifying the important image patches for models than popular methods.
@@ A Visual Set-Sum Task
B Results of the Insertion/Deletion curve with additional models
@@
In Sec. 5.1, we evaluated the proposed and baseline methods using ViT-T [11]. The insertion and deletion curves show that the proposed method provides the most efficient visual explanation. To demonstrate this generalization across different models and architectures, we provide results using both the DeiT-T [26], a ViT architecture, and ResNet-18 [14], a widely used CNN model. For details of the experiment in Deit-T, refer to Sec. 5.1. The insertion curve in Fig. 8(a) again shows that MoXI exhibits a sharper increase compared to the other methods. The deletion curve in Fig. 8(b) also demonstrates that MoXI exhibits a sharper decrease compared to the other methods. Similarly, Fig. 9 exhibits that the results for ResNet-18 are similar to these findings. These results indicate that our method can efficiently and accurately identify the critical patches in the model’s decision-making process.
@@
Figure 8: Results for DeiT-T:(a) Insertion curves. (b) Deletion curves. The curves illustrate the accuracy growth when inserting (deleting) image patches according to the contributions computed by each method.
Figure 9: Results for ResNet-18: (a) Insertion curves. (b) Deletion curves. The curves illustrate the accuracy growth when inserting (deleting) image patches according to the contributions computed by each method.
C Comparison of algorithm complexity and runtime
In this section, we compare the algorithm complexity and runtime for each method.
First, we provide an explanation of the complexity of the algorithm, focusing on the number of forward passes required per image. Let $|N|$ be the number of patches in an image (typically, $|N|=14^{2}$). Grad-CAM needs $1$ forward pass (and $1$ backward pass), Attention rollout needs $1$ forward pass, and Shapley value needs $\mathcal{O}(|N|\cdot 2^{|N|})$ forward passes. MoXI needs $\mathcal{O}(|N|^{2})$ forward passes in the worst case. The number of passes for Shapley value and MoXI is given in Sec. 4, which we will elaborate it again. As defined in Eq. (1), computing the Shapley value for the $i$-th pixel requires $\mathcal{O}(2^{|N|})$ passes due to the $2^{|N|-1}$ possible choices of $S$. leading to $\mathcal{O}(|N|\cdot 2^{|N|})$ passes for an entire image. On the other hand, MoXI needs $O(|N|^{2})$ passes. For example, at the $k$-th step of the greedy insertion, it recruits a new patch from the remaining $|N|-k+1$ patches to maximize the confidence score (i.e., $|N|-k+1$ passes). For $|N|$ steps, it needs $\mathcal{O}(|N|^{2})$ passes in total. Note that this is the worst-case scenario; the algorithm stops when the classification becomes correct, and Fig. 3(a) indicates that more than 90% of evaluation images require less than $0.04N$ steps. In the runtime experiment, the median of the steps was 6 (with std 7.6) and 10 (with std 11.3) for insertion and deletion, respectively. A similar discussion holds for the deletion case.
Furthermore, our method can leverage parallel processing with mini-batches, leading to a linear number of forward passes at the cost of additional memory usage. Specifically, the $k$-th step of MoXI can be done by a single forward pass of $|N|-k+1$ patterns of the insertion from remaining patches. Our implementation is based on this parallelization.
Next, we compare the runtime required for measuring the importance in each method. The comparison is based on the average runtime across 100 images, following the experimental setup described in Sec. 5. For Grad-CAM, Attention rollout, and Shapley value, the runtime represents the duration required to compute the importance of the entire image. In contrast, in the case of MoXI, we separately measure the runtime for pixel insertion until successful classification and for pixel deletion until classification failure. Our experiments were conducted using a machine equipped with a 12-core processor, 64GB RAM, and an NVIDIA RTX 3090.
The runtime for each method is shown in Table 1. This indicates that the runtime for MoXI is approximately $30$ times faster than that for Shapley value. Recalling the results from Fig. 3, MoXI achieves higher accuracy in capturing an important group of patches than Shapley value method does. Therefore, MoXI surpasses Shapley value in both accuracy and runtime. While not as fast as Grad-CAM and Attention rollout, we consider that MoXI meets most use cases of visualization, and the quality is better, as our extensive experiments show.
Table 1: Average runtime 100 ImageNet images [sec] in ViT-T.
Grad-CAM Attention R. Shapley V. MoXI (Ins/Del)
0.15 0.02 17.9 0.60/1.34
D Analysis of effective layers to remove patches
@@
In Sec. 5.1, we consider the absence of players (i.e., pixels/patches) for calculating Shapley values and interactions in the input space. Specifically, the patches are removed after the input embedding layer. Here, we examine the case where several self-attention layers are instead masked. To this end, we utilize a variant of the attention-masking approach used in [8]. Specifically, let the $k$-th layer be our target layer. Then, a large negative value is added to the product of the query and key matrices from $k$-th to the last self-attention layers. Figure 10 displays the insertion curve results when MoXI is applied to various target layers. The experimental setup is the same as in Sec. 5.1. The result demonstrates that MoXI prefers the earlier layers and better pinpoints the important features of images.
@@
E Additional results of visualization
We provide additional visualization results in Fig. 11 and 12. As in Sec. 5.2, the results demonstrated that the patches highlighted by MoXI are smaller than those highlighted by other methods.
We observed that MoXI behaves slightly unstable at the insertion case. Recall that in this case, MoXI appends important patches to an empty set accordingly and terminates when the model gives the correct classification. Empirically, the termination can happen at a very early stage, where the confidence score of the correct class is the largest but still very low. If we continue to patch, the model prediction can fluctuate among several classes. Note this does not cause a big problem in most cases; all the insertion curves in this paper consistently show a monotonic increase of classification accuracy with the increase in insertion rate. If needed, one can introduce a minimum confidence score $\tau\in[0,1]$ and terminate the insertion when the confidence score exceeds this threshold with the correct classification. We include this hyperparameter in our official implementation of MoXI.